OpenStudy (anonymous):

Verify this:

OpenStudy (anonymous):

\[(\tan \theta + \sec \theta-1) /(\tan \theta - \sec \theta + 1) = 1 + \sin \theta / \cos \theta\]

OpenStudy (anonymous):

@gorv

OpenStudy (paxpolaris):

rewrite the LHS using only sin and cos

OpenStudy (anonymous):

thats what I'm trying to do and I can't figure it out

OpenStudy (anonymous):

like sin/cos? and 1/cos?

OpenStudy (paxpolaris):

right

OpenStudy (anonymous):

ok then what?

OpenStudy (paxpolaris):

\[LHS\\={\tan \theta +\sec \theta -1 \over \tan \theta -\sec \theta+1}\\ \large=\color{blue}{{\sin \theta \over \cos \theta} +{1 \over \cos \theta}-1 \over {\sin \theta \over \cos \theta} -{1 \over \cos \theta}+1 }\]

OpenStudy (anonymous):

yes that is what i got

OpenStudy (anonymous):

then do yo do common denominators for both sides

OpenStudy (anonymous):

you*

OpenStudy (paxpolaris):

how about we next multiply the numerator and denom by cos(theta)

OpenStudy (anonymous):

i meant both top and bottom

OpenStudy (anonymous):

ok so you get \[\sin \theta - \cos \theta +1 / \sin \theta + \cos \theta - 1\]

OpenStudy (anonymous):

you don't have to spend so much time writing the equation you can just explain it to me

OpenStudy (paxpolaris):

\[={\sin \theta+1- \cos \theta \over \sin \theta-1+ \cos \theta}\\={\sin \theta- (\cos \theta-1) \over \sin \theta+ (\cos \theta-1)}\] maybe multiplying the numerator to both the numerator and denom would help next

OpenStudy (anonymous):

what would the numerator be to multiply, that whole expression?\[\sin \theta - (\cos \theta - 1)\]

OpenStudy (paxpolaris):

yes ...i am thinking that would help

OpenStudy (anonymous):

i don't know what the result of multiplying the numerator would be

OpenStudy (anonymous):

ok so i got \[\sin ^2 \theta - 2\sin \theta \cos \theta + 2 \sin \theta + ( \cos \theta - 1) ^2 / \sin ^2 \theta - ( \cos \theta - 1) ^2\]

OpenStudy (paxpolaris):

\[\color{blue}{\sin^2\theta} -2\sin \theta(\cos \theta -1)\color{blue}{+\cos^2\theta}-2\cos \theta+1 \over \color{blue}{\sin^2\theta} - \cos^2 \theta+2\cos \theta \color{blue}{-1}\]

OpenStudy (anonymous):

ok thats what i got I guess and then what

OpenStudy (paxpolaris):

replace what i marked in blue sin^2+cos^2=1 sin^2-1= -cos^2

OpenStudy (anonymous):

ok and then what

OpenStudy (paxpolaris):

\[=-{2\sin \theta(\cos \theta-1)-2\cos+2 \over -2\cos^2\theta+2\cos \theta}\] looks like numerator and denom have something in common ... reduce

OpenStudy (paxpolaris):

**\[={\color{red}-2\sin \theta(\cos \theta-1)-2\cos+2 \over -2\cos^2\theta+2\cos \theta}\]

OpenStudy (anonymous):

ok what is the reduced form?

OpenStudy (paxpolaris):

\[={-2\left( \cos \theta-1 \right)\left[ \sin \theta+1 \right] \over -2\left( \cos \theta-1 \right)\cos \theta}\]

OpenStudy (anonymous):

thanks